# Fine-tuning Whisper on Low-Resource Languages for Real-World Applications

### v1

@@ Abstract
This paper presents a new approach to fine-tuning OpenAI’s Whisper model for low-resource languages by introducing a novel data generation method that converts sentence-level data into a long-form corpus, using Swiss German as a case study. Non-sentence-level data, which could improve the performance of long-form audio, is difficult to obtain and often restricted by copyright laws. Our method bridges this gap by transforming more accessible sentence-level data into a format that preserves the model’s ability to handle long-form audio and perform segmentation without requiring non-sentence-level data. Our data generation process improves performance in several real-world applications and leads to the development of a new state-of-the-art speech-to-text (STT) model for Swiss German. We compare our model with a non-fine-tuned Whisper and our previous state-of-the-art Swiss German STT models, where our new model achieves higher BLEU scores. Our results also indicate that the proposed method is adaptable to other low-resource languages, supported by written guidance and code that allows the creation of fine-tuned Whisper models, which keep segmentation capabilities and allow the transcription of longer audio files using only sentence-level data with high quality.
keywords:
speech recognition, Whisper, fine-tune, low-resource, speech translation, Swiss German
@@ 1 Introduction
Swiss German refers to the dialects spoken in the German-speaking regions of Switzerland. Due to the limited number of speakers, linguistic resources are scarce, also because Swiss German exists only as a spoken language, without any formal grammar or standardized written form. As a result, Swiss German STT systems are typically formulated as speech translation tasks, where Swiss German audio is transcribed into standard German text [1, 2, 3].
The Whisper models [4] developed by OpenAI are trained on a large-scale corpus of audio recordings and corresponding transcriptions obtained by web crawling. The dataset used for the multilingual version of Whisper includes samples from almost 100 different languages. After English and Chinese, German represents the third-largest part of the dataset, with 13’344 hours. The unexpectedly high transcription quality of Whisper for Swiss German audio and video (see Table 4) and observed distinct hallucinations (discussed in Section 5) prove the presence of Swiss German audio in the original training dataset. While the model performs remarkably well for Swiss German, there is still considerable room for improvement for practical applications that require higher transcription quality, such as judicial interrogation transcripts or medical diagnosis and treatment orders.
@@
Table 1: Comparison of a rapidly spoken Swiss-German saying [5], transcribed by Whisper Large-v2. The fine-tuned model fails to predict timestamps and performs worse than the original Large-v2.
@@ 2 Related work
Despite the challenge of scarce data resources, recent advancements in speech translation for Swiss German have been substantial, driven in part by recent collections of high-quality sentence-level datasets from Swiss parliaments minutes and crowdsourcing initiatives such as SwissDial [6], the Swiss Parliaments Corpus [1], SDS-200 [2], and STT4SG-350 [3].
Prior to the release of Whisper, a commonly applied foundation model for building ASR systems was XLS-R [7]. XLS-R is based on the wav2vec 2.0 architecture [8] and was pre-trained on 436K hours of speech data in 128 languages. Previous research on Swiss German speech recognition has therefore often used the pre-trained XLS-R backbone [2, 3, 9, 10].
There are many papers and blog posts on the fine-tuning of Whisper [11, 12, 13, 14, 15, 16, 17, 18, 19] and some focused on low-resource settings [20, 21, 22, 23]. The problem of language forgetting is also discussed extensively, and it is shown that fine-tuning on a new language yields the best performance for the new language, but degrades the capabilities on existing languages [24]. However, it is rare to find papers that explicitly address the problem of fine-tuning Whisper for transcription of longer audios. Many papers also do not evaluate or discuss the segmentation capabilities after fine-tuning. Finally, it is uncommon to see evaluations of fine-tuned Whisper on out-of-distribution datasets.
@@ 3.2 Training Details
Using gradient check-pointing [25], gradient accumulation, and an 8-bit optimizer [26], we achieve an effective batch size of 256 on a single NVIDIA A100 40 GB GPU. Gradient check-pointing is applied to both the encoder and decoder, with 16 gradient accumulation steps and a per-step batch size of 16. Additionally, we apply stochastic depth [27] to the encoder and decoder blocks. This setup results in a mixed-precision training run that takes about 42 hours.
We use a learning rate scheduler with a linearly increasing warm-up phase followed by a linear decay to zero, as described in the original Whisper training procedure [4]. During training, each sample has a 50% chance of containing timestamps and a 50% chance of containing prompts, mentioned in a comment on the OpenAI Whisper repository  [^3] https://github.com/openai/Whisper/discussions/838.
@@
Following the improved training procedure of Whisper Large-v2, we apply SpecAugment [28] during training with the same parameters as in [4], summarized in Table 2.
@@ 3.3 Train, Validation and Test Data
For our training data, we use the Swiss German sentence-level datasets [1, 2, 3] with the predefined train and validation sets. For the train and validation split, unless mentioned otherwise, we use our data-generation pipeline explained in section 3.1. As additional training and validation data, we use Swiss Broadcasting Corporation (SRG) shows, pseudo-labeled (PL) by transcription with Whisper Large-v2. We selected 17 TV series, in which Swiss German is spoken.
@@
In Figure 3, we analyze the relationship between the BLEU score [29] on the STT4SG-350 test set and the amount of training data used for fine-tuning. For training, we used the 502 hours long-form corpus consisting of SDS-200, STT4SG-350 and SPC, but we do not include the pseudo-labeled data to show what can be expected from high-quality labeled data. The models were trained using hierarchically nested datasets, each partition holding approximately 20% of the training data. Training was run until the word error rate (WER) on the validation set showed no more improvement. Once the training had stabilized, the best performing model on WER was selected.
@@ 4.1 Evaluation
In all experiments, the computed metrics are derived from the WhisperX [^4] https://github.com/m-bain/whisperX versions of the models using fp16 precision. The beam size is configured to 5, VAD is enabled and the language tag is DE. For the sentence-level datasets, we report the WER or the BLEU metric or both, if the layout allows it. For the long-form test dataset we replace the BLEU metric with SubER [30]; this allows us to incorporate a metric that also assesses the segmentation quality. Before calculating the metrics, all sentences are transformed to lowercase, and punctuation is removed. For the BLEU metric we specifically use sacreBLEU [31] with default parameters.
@@ 4.2 Base Results
Table 4: Performance of the original Whisper models on various test sets. SubER is shown on the long-form SRG data.
@@ 4.4 Overcome Shortcomings
As part of our methodology to mitigate language forgetting in the final model training, we use the Mozilla Common Voice 13 German dataset [32]. Using the data generation pipeline described in Section 3.1, we curated a subset of 15,000 samples, each lasting 30 seconds, resulting in 125 hours of additional training data. The train, validation, and test set splits were taken as defined by Mozilla Common Voice Version 13.
@@ 4.5 Dialect Comparison
The results in Table 7 show large differences in the performance of the original Large-v2 model across the different dialect regions. In contrast, our fine-tuned SOTA model exhibits improved WER over a much narrower range across all dialects. For the reader, the improvements over our older models [9] are also shown.
Table 7: WER for Swiss German dialects on the STT4SG-350 test set for selected models; XLS-R and TF are older models published in [9].
@@ 5 Conclusions
In addition, we have highlighted that while the model may improve on data from the same distribution as the training data, in reality the model gets worse, as shown in table 5. This underscores the importance of creating or acquiring evaluation datasets that closely mimic the intended deployment environment, ensuring the ASR system’s robustness and usefulness.
@@
Additionally, we observed distinct hallucinations of Whisper Large-v2 mentioning Swiss subtitling companies, such as being able to reliably trigger Whisper to transcribe ”Untertitel von SWISS TXT” - a watermark of SWISS TXT that is only present in the subtitle files, never in the audio - when asked to transcribe the title music of the SRF Meteo show or when music is being played in ”SRF bi de Lüt”.
@@ 6 Future work
As we have extensively analyzed and evaluated different methods to generate long-form data from sentence-level data, the combination of the data generation and methods to avoid catastrophic forgetting, as presented by [24] by using Elastic Weight Consolidation [33], could be a next research topic. Another potential next step involves diversifying the data sources by augmenting the pseudo-labeled datasets with additional real-world data, including a broader range of TV programs, varied conversational contexts, and noisy environments. This expansion aims to enhance the robustness and generalization capabilities of the models. Notably, preliminary experiments indicate that a fine-tuned Whisper Large-v3 model performs particularly well on conversational speech, highlighting its potential superiority in this context and emphasizing the need for a large corpus of freely spoken dialogues.

### v2

@@ Abstract
This paper presents a new approach to fine-tuning OpenAI’s Whisper model for low-resource languages by introducing a novel data generation method that converts sentence-level data into a long-form corpus, using Swiss German as a case study. Non-sentence-level data, which could improve the performance of long-form audio, is difficult to obtain and often restricted by copyright laws. Our method bridges this gap by transforming more accessible sentence-level data into a format that preserves the model’s ability to handle long-form audio and perform segmentation without requiring non-sentence-level data. Our data generation process improves performance in several real-world applications and leads to the development of a new state-of-the-art speech-to-text (STT) model for Swiss German. We compare our model with a non-fine-tuned Whisper and previous state-of-the-art Swiss German STT models, where our new model achieves higher BLEU scores. Our results also indicate that the proposed method is adaptable to other low-resource languages, supported by written guidance and code that allows the creation of fine-tuned Whisper models, which keep segmentation capabilities and allow the transcription of longer audio files using only sentence-level data with high quality.
Fine-tuning Whisper on Low-Resource Languages for Real-World Applications
Vincenzo Timmel1, Claudio Paonessa2, Manfred Vogel1, Daniel Perruchoud1, Reza Kakooe1 1University of Applied Sciences and Arts Northwestern Switzerland 2Noxenum.io {vincenzo.timmel, manfred.vogel, daniel.perruchoud, reza.kakooee}@fhnw.ch, claudio.paonessa@noxenum.io
@@ 1 Introduction
Swiss German refers to the dialects spoken in the German-speaking regions of Switzerland. Due to the limited number of speakers, linguistic resources are scarce, also because Swiss German exists only as a spoken language, without any formal grammar or standardized written form. As a result, Swiss German STT systems are typically formulated as speech translation tasks, where Swiss German audio is transcribed into standard German text Plüss et al. (2021, 2022, 2023).
The Whisper models Radford et al. (2022) developed by OpenAI are trained on a large-scale corpus of audio recordings and corresponding transcriptions obtained by web crawling. The dataset used for the multilingual version of Whisper includes samples from almost 100 different languages. After English and Chinese, German represents the third-largest part of the dataset, with 13’344 hours. The unexpectedly high transcription quality of Whisper for Swiss German audio and video (see Table 4) and observed distinct hallucinations (discussed in Section 5) prove the presence of Swiss German audio in the original training dataset. While the model performs remarkably well for Swiss German, there is still considerable room for improvement for practical applications that require higher transcription quality, such as judicial interrogation transcripts or medical diagnosis and treatment orders.
@@
Table 1: Comparison of a rapidly spoken Swiss-German saying Straub (2024), transcribed by Whisper Large-v2. The fine-tuned model fails to predict timestamps and performs worse than the original Large-v2.
@@ 2 Related work
Despite the challenge of scarce data resources, recent advancements in speech translation for Swiss German have been substantial, driven in part by recent collections of high-quality sentence-level datasets from Swiss parliaments minutes and crowdsourcing initiatives such as SwissDial Dogan-Schönberger et al. (2021), the Swiss Parliaments Corpus Plüss et al. (2021), SDS-200 Plüss et al. (2022), and STT4SG-350 Plüss et al. (2023).
Prior to the release of Whisper, a commonly applied foundation model for building ASR systems was XLS-R Babu et al. (2021). XLS-R is based on the wav2vec 2.0 architecture Baevski et al. (2020) and was pre-trained on 436K hours of speech data in 128 languages. Previous research on Swiss German speech recognition has therefore often used the pre-trained XLS-R backbone Plüss et al. (2022, 2023); Schraner et al. (2022); Paonessa et al. (2023).
There are many papers and blog posts on the fine-tuning of Whisper Gandhi (2022); Singh (2023); de Bruin (2023); OpenAI (2023); Ma et al. (2023); Shamsian et al. (2024); Do et al. (2023); Ferraz et al. (2023); Sicard et al. (2023) and some focused on low-resource settings Piñeiro-Martín et al. (2024); Hsu et al. (2024); Liu et al. (2024); Pillai et al. (2024). The problem of language forgetting is also discussed extensively, and it is shown that fine-tuning on a new language yields the best performance for the new language, but degrades the capabilities on existing languages Qian et al. (2024). However, it is rare to find papers that explicitly address the problem of fine-tuning Whisper for transcription of longer audios. Many papers also do not evaluate or discuss the segmentation capabilities after fine-tuning. Finally, it is uncommon to see evaluations of fine-tuned Whisper on out-of-distribution datasets.
@@ 3.2 Training Details
Using gradient check-pointing Sohoni et al. (2022), gradient accumulation, and an 8-bit optimizer Dettmers et al. (2021), we achieve an effective batch size of 256 on a single NVIDIA A100 40 GB GPU. Gradient check-pointing is applied to both the encoder and decoder, with 16 gradient accumulation steps and a per-step batch size of 16. Additionally, we apply stochastic depth Huang et al. (2016) to the encoder and decoder blocks. This setup results in a mixed-precision training run that takes about 42 hours.
We use a learning rate scheduler with a linearly increasing warm-up phase followed by a linear decay to zero, as described in the original Whisper training procedure Radford et al. (2022). During training, each sample has a 50% chance of containing timestamps and a 50% chance of containing prompts, mentioned in a comment on the OpenAI Whisper repository  [^3] https://github.com/openai/Whisper/discussions/838.
@@
Following the improved training procedure of Whisper Large-v2, we apply SpecAugment Park et al. (2019) during training with the same parameters as in Radford et al. (2022), summarized in Table 2.
@@ 3.3 Train, Validation and Test Data
For our training data, we use the Swiss German sentence-level datasets Plüss et al. (2021, 2022, 2023) with the predefined train and validation sets. For the train and validation split, unless mentioned otherwise, we use our data-generation pipeline explained in section 3.1. As additional training and validation data, we use Swiss Broadcasting Corporation (SRG) shows, pseudo-labeled (PL) by transcription with Whisper Large-v2. We selected 17 TV series, in which Swiss German is spoken.
@@
In Figure 3, we analyze the relationship between the BLEU score Papineni et al. (2002) on the STT4SG-350 test set and the amount of training data used for fine-tuning. For training, we used the 502 hours long-form corpus consisting of SDS-200, STT4SG-350 and SPC, but we do not include the pseudo-labeled data to show what can be expected from high-quality labeled data. The models were trained using hierarchically nested datasets, each partition holding approximately 20% of the training data. Training was run until the word error rate (WER) on the validation set showed no more improvement. Once the training had stabilized, the best performing model on WER was selected.
@@ 4.1 Evaluation
In all experiments, the computed metrics are derived from the WhisperX [^4] https://github.com/m-bain/whisperX versions of the models using fp16 precision. The beam size is configured to 5, VAD is enabled and the language tag is DE. For the sentence-level datasets, we report the WER or the BLEU metric or both, if the layout allows it. For the long-form test dataset we replace the BLEU metric with SubER Wilken et al. (2022); this allows us to incorporate a metric that also assesses the segmentation quality. Before calculating the metrics, all sentences are transformed to lowercase, and punctuation is removed. For the BLEU metric we specifically use sacreBLEU Post (2018) with default parameters.
@@ 4.2 Base Results
Table 4: Performance of the original Whisper models on various test sets. SubER is shown on the long-form SRG data. *
@@ 4.4 Overcome Shortcomings
As part of our methodology to mitigate language forgetting in the final model training, we use the Mozilla Common Voice 13 German dataset Ardila et al. (2020). Using the data generation pipeline described in Section 3.1, we curated a subset of 15,000 samples, each lasting 30 seconds, resulting in 125 hours of additional training data. The train, validation, and test set splits were taken as defined by Mozilla Common Voice Version 13.
@@ 4.5 Dialect Comparison
The results in Table 7 show large differences in the performance of the original Large-v2 model across the different dialect regions. In contrast, our fine-tuned SOTA model exhibits improved WER over a much narrower range across all dialects. For the reader, the improvements over other Swiss-German ASR models Schraner et al. (2022) are also shown.
Table 7: WER for Swiss German dialects on the STT4SG-350 test set for selected models; XLS-R and TF are older models published in Schraner et al. (2022).
@@ 5 Conclusions
In addition, we have highlighted that while the model may improve on data from the same distribution as the training data, in reality the model preforms worse on out-of-distribution data, as shown in table 5. This underscores the importance of creating or acquiring evaluation datasets that closely mimic the intended deployment environment, ensuring the ASR system’s robustness and usefulness in real-world applications.
@@
Additionally, we observed distinct hallucinations of Whisper Large-v2 mentioning Swiss subtitling companies, such as being able to reliably trigger Whisper to transcribe "Untertitel von SWISS TXT" - a watermark of SWISS TXT that is only present in the subtitle files, never in the audio - when asked to transcribe the title music of the SRF Meteo show or when music is being played in "SRF bi de Lüt".
@@ 6 Future work
As we have extensively analyzed and evaluated different methods to generate long-form data from sentence-level data, the combination of the data generation and methods to avoid catastrophic forgetting, as presented by Qian et al. (2024) by using Elastic Weight Consolidation Kirkpatrick et al. (2017), could be a next research topic. Another potential next step involves diversifying the data sources by augmenting the pseudo-labeled datasets with additional real-world data, including a broader range of TV programs, varied conversational contexts, and noisy environments. This expansion aims to enhance the robustness and generalization capabilities of the models. Notably, preliminary experiments indicate that a fine-tuned Whisper Large-v3 model performs particularly well on conversational speech, highlighting its potential superiority in this context and emphasizing the need for a large corpus of freely spoken dialogues.